OpenStudy (anonymous):

Simplify the expression.(7 – c)(–1) A.) 7+c B.) 7-c C.)-7-c D.)-7+c

OpenStudy (shaik0124):

d

OpenStudy (shaik0124):

-7-(-c) -7+c

OpenStudy (anonymous):

(7-c)(-1) = -1*7 - (c)(-1) = -7 +c

OpenStudy (mathstudent55):

If it'll make it look simpler, rewrite it with the -1 on the left side. Then use the distributive property. a(b + c) = ab + ac